OpenStudy (yanasidlinskiy):

a car traveling at 27m/s uniformly accelerates to rest over distance of 60.8m. What is the deceleration rate? How long did it take to stop?

jigglypuff314 (jigglypuff314):

Here are the equations we should use, they are commonly known as "Kinematic Equations" 

jigglypuff314 (jigglypuff314):

oh... it got cut off ;-; http://www.physicsclassroom.com/Class/1DKin/U1L6a1.gif

OpenStudy (yanasidlinskiy):

It's ok. I have those.

jigglypuff314 (jigglypuff314):

First step is to list the information we do have so \(\large v_i = 27 \frac{m}{s}\) \(\large v_f = 0 \frac{m}{s}\) \(d = 60.8m\) and we have two unknowns a = ? t = ? make sense so far? :)

OpenStudy (yanasidlinskiy):

Yes.

OpenStudy (yanasidlinskiy):

Hold on.

jigglypuff314 (jigglypuff314):

Good then to find \(a\) we look for the equation that doesn't use \(t\)

OpenStudy (yanasidlinskiy):

Thts what it is. Ok and we'll use the top right correct?

OpenStudy (yanasidlinskiy):

Like I don't get wht I'm supposed to fill in the bottom blanks(graph)

jigglypuff314 (jigglypuff314):

so for each problem you fill out the information we know (like I did above) and use the equations given at the top to find the missing variables

OpenStudy (yanasidlinskiy):

Okay. Can we work the first one together so I kinda know how to do the rest?

jigglypuff314 (jigglypuff314):

A car uniformly accelerates from rest to a final speed of 88 mph (39.2 m/s) in 4 seconds. What is the acceleration rate of the car? How far did it travel? ^ this one right?

OpenStudy (yanasidlinskiy):

Yes.

OpenStudy (yanasidlinskiy):

I had 78.4 in the first box then v(m/s) = 39.2 a(m/s^2) = 9.8 t(sec) = 4 I don't know the answer for V0(m/s)

jigglypuff314 (jigglypuff314):

\(v_0\) means the starting speed for this, the car starts "from rest" which means that it's starting speed is 0 m/s

OpenStudy (yanasidlinskiy):

So 0?

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (yanasidlinskiy):

Okay. Now the second one lol

jigglypuff314 (jigglypuff314):

A car traveling at 27 m/s uniformly accelerates to rest over a distance of 60.8m. What is the deceleration rate? How long did it take to stop? so the starting speed is 27 m/s the ending speed is 0 (because it "accelerates to rest") distance is 60.8 and we would have to solve for the acceleration and time

OpenStudy (yanasidlinskiy):

So 60.8 goes in the first box right?

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (yanasidlinskiy):

27 goes in the second?

jigglypuff314 (jigglypuff314):

correct ^_^

OpenStudy (yanasidlinskiy):

Okay so wht would I put down for the v?

jigglypuff314 (jigglypuff314):

it says "accelerates to rest" "rest" means a speed/velocity of 0

OpenStudy (yanasidlinskiy):

gotcha

OpenStudy (yanasidlinskiy):

Okay this part confuses me cuz idk which formulas to use for which questions

jigglypuff314 (jigglypuff314):

We have two missing things acceleration and time if we want to find acceleration, we use the equation that is missing time if we want to find time, we use the equation that is missing acceleration

OpenStudy (yanasidlinskiy):

I still don't get it. All the problems have either one or the other...

jigglypuff314 (jigglypuff314):

Let's try finding acceleration first together then? :) we would use the equation that is missing time

OpenStudy (yanasidlinskiy):

Okay

jigglypuff314 (jigglypuff314):

\((v)^2 = (v_0)^2 + 2ad\) this equation ^_^

jigglypuff314 (jigglypuff314):

then we plug in what we already know

jigglypuff314 (jigglypuff314):

so plug in and what would we get? :)

OpenStudy (yanasidlinskiy):

0^2= 27^2+2(idk)(60.8) Is tht right?

jigglypuff314 (jigglypuff314):

yes :) then we try to solve for (idk) like in algebra

OpenStudy (yanasidlinskiy):

Hahha ok lol

OpenStudy (yanasidlinskiy):

I got 850.6

jigglypuff314 (jigglypuff314):

ehhh can't do it like that think of it like this 0= 27^2+2(60.8)x and "solve for x"

OpenStudy (yanasidlinskiy):

Im getting the same thing

jigglypuff314 (jigglypuff314):

0 = 27^2+2(60.8)x -(27^2) -(27^2) minus (27^2) from both sides ----------------------------------- -(27^2) = (2*60.8)x /(2*60.8) /(2*60.8) divide both sides by (2*60.8) ----------------------------------- \( \large x = \frac{-(27^2)}{2*60.8}\)

OpenStudy (yanasidlinskiy):

Ohhhhhhhh ok hold up

OpenStudy (yanasidlinskiy):

I was using a calculator

jigglypuff314 (jigglypuff314):

you can't use calculators when there are variables

OpenStudy (yanasidlinskiy):

Yea ik. i wanted to get this over with lol

jigglypuff314 (jigglypuff314):

that ^ would get you your acceleration

OpenStudy (yanasidlinskiy):

So -6.0? I rounded it off from -5.99

jigglypuff314 (jigglypuff314):

yup ^_^

OpenStudy (yanasidlinskiy):

Okay so now time. Would I use this?: v=v0+at?

jigglypuff314 (jigglypuff314):

use the one without acceleration (because that was rounded and inaccurate) d = (1/2)*(v + v_0)*t

OpenStudy (yanasidlinskiy):

Okay

jigglypuff314 (jigglypuff314):

\( \large d = \frac{v+v_0}{2} * t \) d = 60.8 v = 0 v0 = 27 and we solve for t

OpenStudy (yanasidlinskiy):

0+27/2 *T 13.5?

jigglypuff314 (jigglypuff314):

60.8 = 13.5 t now divide both sides by 13.5

OpenStudy (yanasidlinskiy):

4.5

jigglypuff314 (jigglypuff314):

there you go :)

OpenStudy (yanasidlinskiy):

Okay. I'm still iffy on these problems lol

OpenStudy (yanasidlinskiy):

Can we do a few more if you don't mind?

jigglypuff314 (jigglypuff314):

Try some yourself, it's basically Algebra 1 problems...

OpenStudy (yanasidlinskiy):

Ok

jigglypuff314 (jigglypuff314):

problems like these hide pretty well as math problems (so you can ask it in the math section) ^_^"

OpenStudy (yanasidlinskiy):

Okay. Thanks.

OpenStudy (yanasidlinskiy):

I got 5.76 for number 3. Is that right?